OpenStudy (31356):

A small plane can carry 800 pounds of baggage, distributed between two storage compartments. On one flight the plane is fully loaded, with 98 pounds more baggage in one compartment than the other. How much is stored in each compartment?

OpenStudy (31356):

I got x+y=800 And y=x+98 what am I supposed to do? @hero and @douglaswinslowcooper ?

OpenStudy (31356):

And @dumbcow?

OpenStudy (anonymous):

good start y+x = 800 y-x = 98 add these to get rid of x , 2y = 898 and solve for y, then for x

OpenStudy (31356):

Okay thanks :)(:

OpenStudy (anonymous):

You are quite welcome.